# Navier-Stokes Equations with Fractional Dissipation and Associated Doubly Stochastic Yule Cascades

### v1

@@ 3 $(d,\gamma)$-DSY cascade
The quadratic form $a\odot_{\xi}b$ defined in (2.8) satisfies the following estimate:
@@
The proof of (4.1) is elementary and can be found in [dascaliuc2019jan, Lem. 2.1].
To establish the integrability of the solution process $\mathbf{X}$ defined (2.14) it is natural to use the majorization (4.1) of the quadratic form $a\odot_{\xi}b$. In so doing, we establish a natural relation with equation (MS), which will be referred in this paper as the (fractional) Montgomery-Smith equation. The author of [smith] introduce (MS) for the case $d=3,\gamma=1$, as a tool of explore the limitation of methods of functional analysis in establishing the well possednes of the Navier-Stokes equations. In the context of $(d,\gamma)$-DSY cascades, the Montgomery-Smith equation plays two roles. The first role is as the governing equation for the probability of no explosion
$\rho(\xi,t)=\mathbb{P}_{\xi}(S>t)$
@@
For the first, recall that in Section 3 of the $(d,\gamma)$-DSY cascade, the distribution of the wave vector $W_{v1}$ given $W_{v}=\xi$ is $H(\eta|\xi)=h(\eta)h(\xi-\eta)/(|\xi|^{2\gamma-1}h(\xi))$ where $h$, given by (2.10), is a positive function satisfying the fundamental relation (2.11) $h*h(\xi)=|\xi|^{2\gamma-1}h(\xi)$. We then have,
@@
Consequently, $\rho$ is continuous in time.
@@
This follows similar lines as the proof of Section 2. Note that $\mathscr{X}(\xi,t)=\mathbbm{1}_{S_{\xi}>t}$ satisfies
$\mathscr{X}(\xi,t)=\left\{\begin{array}[]{ll}1,&Y_{\emptyset}\geq t,\\ \mathscr{X}^{(1)}(\eta,t-Y_{\emptyset})\mathscr{X}^{(2)}(\xi-\eta,t-Y_{\emptyset}),&Y_{\emptyset}<t,\end{array}\right$. (4.3)
where $\mathscr{X}^{(1)}$ and $\mathscr{X}^{(2)}$ are two conditionally independent copies on $\mathscr{X}$. Indeed, this follows from the inclusion $Y_{\emptyset}\geq t]\subset[S>t]$ and from the inherited nature of the non-explosion event: ${\mathbbm{1}_{S>t}}\mathbbm{1}_{Y_{\emptyset}<t}=\mathbbm{1}_{Y_{\emptyset}<t}\left(\mathbbm{1}_{{{S}_{{{W}_{1}}}}>t-Y_{\emptyset}}+\mathbbm{1}_{{S_{{{W}_{2}}}}>t-Y_{\emptyset}}\right)$. Since $|\bar{\mathbf{X}}|\leq 1$, we can take expectation in (4.3) to obtain that $\rho=\mathbb{E}\bar{\mathbf{X}}$ satisfies (4.2). The continuity of $\rho$ follows immediately from (4.2) since $\int_{\mathbb{R}^{3}}\rho(\eta,t-s)\rho(\xi-\eta,t-s)H(\eta|\xi)d\eta\leq 1.$
Note that the constant function $1$ is a trivial solution of (4.2). Consequently, determining whether stochastic explosion occurs (i.e. $\rho(\xi,t)<1$ for some $(\xi,t)$) is equivalent to determining whether (4.2) has other solutions besides the trivial solution. This is indeed the case, since as shown in Section 3 there exists values of $(d,\gamma)$ such that the resulting DSY cascade is explosive.
Generalizing the equation introduced in [smith], consider the fractional Montgomery-Smith equation in $d$ dimensions given by
@@
with initial data $v_{0}(\xi)$. Here, $v(\xi,t)$ and $v_{0}(\xi)=v(\xi,0)$ are scalar valued functions. Note that (MS) has the same scaling property as the fractional Navier-Stokes equations (NSE)d,γ. The Fourier transform of (MS) is given, with $c_{0}=(2\pi)^{-d/2}$ by
@@
In what follows it will be necessary to introduced the normalized Fourier Montgomery-Smith equation given by
@@
$\mathbb{E}_{\xi}[|\mathbf{X}|\mathbbm{1}_{S>t}]\leq 2\mathbb{E}_{\xi}\bar{\mathbf{X}}<\infty$
@@
According to Section 4(c), if (FMS) has a mild solution on $(0,\tau)$ then $\mathbb{E}_{\xi}[|{\bf{X}}|\mathbbm{1}_{S>t}]<\infty$ a.e. on $\mathbb{R}^{d}\times(0,\tau)$ which infers the lower bound on the critical time $t_{c}\geq\tau$. The following result (due to Ferreira and Villamizar-Roa [ferreira]) asserts the local-in-time existence and uniqueness of mild solutions to (FMS) for initial data $v_{0}$ in the pseudomeasure-type space $PM^{a}$ defined by (2.3).
@@
Note that the proof of Theorem 1.2, 1.3, and Remark 1.5(b) of [ferreira] for the special case $a=b$ and $\theta=1$ does not use any special structure of the $\odot$-product other than the fact that $|a\odot_{\xi}b|\leq|a||b|$. Thus, the results hold for the fractional Montgomery-Smith equation. ∎
@@
We begin by obtaining an expression of the solution operator that involves a product of scalar quantities determined in terms of the initial data, and the angles between the labels of the branching process.
@@
Let $N(\xi,t)=|\hskip-3.61371pt\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)|$. We will prove by induction on $N(\xi,t)$. If $N(\xi,t)=0$ then $\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)=\emptyset$, $\partial V(\xi,t)=\{\theta\}$ and
@@
Thus, (5.1) holds for the base case $N(\xi,t)=0$. Suppose that (5.1) holds for any $N(\xi,t)\leq n$, for some $n\geq 0$. For $N(\xi,t)=n+1\geq 1$, we have $Y_{\theta}<t$. Thus,
$\mathbf{X}(\xi,t)=\mathbf{X}^{(1)}(W_{1},t-Y_{\theta})\odot_{\xi}\mathbf{X}^{(2)}(W_{2},t-Y_{\theta}).$ (5.2)
@@
$\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)$ $=$ $\{\theta\}\cup\stackrel{{\scriptstyle\circ}}{{V}}^{(1)}(W_{1},t-Y_{\theta})\cup\stackrel{{\scriptstyle\circ}}{{V}}^{(2)}(W_{2},t-Y_{\theta})$ (5.3)
$\partial V(\xi,t)$ $=$ $\partial V^{(1)}(W_{1},t-Y_{\theta})\cup\partial V^{(2)}(W_{2},t-Y_{\theta})$ (5.4)
where $\stackrel{{\scriptstyle\circ}}{{V}}^{(k)}$ and $\partial V^{(k)}$ are the set of internal nodes and the set of leaves of the subtree rooted at $W_{k}$. Since $N^{(1)}(W_{1},t-Y_{\theta})+N^{(2)}(W_{2},t-Y_{\theta})=N(\xi,t)-1=n$, one deduces by the induction hypothesis that for $k\in\{1,2\}$,
$\mathbf{X}^{(k)}(W_{k},t-Y_{\theta})$ $=$ $\left(\frac{i}{2}\right)^{N^{(k)}(W_{k},t-Y_{\theta})}\cdot$
$\left(\prod_{v\in\stackrel{{\scriptstyle\circ}}{{V}}^{(k)}(W_{k},t-Y_{\theta})}\textup{sign}(W_{v}\times W_{v1})\sin(\theta_{W_{v},W_{v1}}-\theta_{W_{v},W_{v2}})\right)\cdot$
$\left(\prod_{v\in\partial V^{(k)}(W_{k},t-Y_{\theta})}\chi_{0}(W_{v})\cdot e_{W_{v}^{\perp}}\right)e_{W_{k}^{\perp}}$
@@
Let $\tilde{\mathbf{X}}(\xi,t)=\phi_{\xi,t}(\{\tilde{\Sigma}_{v}\}_{v\in\mathbb{T}},\{\tilde{\Theta}_{v}\}_{v\in\mathbb{T}})$ where $\{\tilde{\Sigma}_{v}\}_{v\in\mathbb{T}}$ and $\{\tilde{\Theta}_{v}\}_{v\in\mathbb{T}}$ are given by (5.6)-(5.9). Suppose the function $\chi_{0}(\eta)\cdot e_{\eta^{\perp}}$ is radially symmetric, i.e. only depending on $|\eta|$. Then
@@
The radial symmetry of the initial data and the invariance of the DSY process leads to the following consequence.
@@
Suppose $u_{0}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ is a vortex. Then (NSE)2,γ has a global solution $u$ whose Fourier transform is given by $\hat{u}(\xi,t)=\hat{u}_{0}(\xi)e^{-|\xi|^{2\gamma}t}$.
@@
${\mathbf{X}}^{*}=\hat{u}_{0}(\xi)\mathbbm{1}[T_{\emptyset}>t].$
@@
For each $r>0$, let $D_{r}=\{\xi\in\mathbb{R}^{2}:|\xi|<r\}$. By the continuity of the translation operator in $L^{1}(\mathbb{R}^{2})$ (see e.g. [rudin, Exer. 12, p. 132]),
@@
Thus, $q(\tau)\geq\limsup_{t\to\tau^{-}}p(t)=\infty$. This is a contradiction. ∎ The fact that the solution process $\mathbf{X}(\xi,t)$ fails to be in integrable after finite time for sufficiently large initial condition does not necessarily imply that the mild solution $u$ to (NSE)d,γ blows up after finite time. It simply implies that the mean-flow representation of the scaled Fourier transform $\chi=c_{0}\frac{\hat{u}}{h}$ ceases to be valid after finite time. For example, if the initial velocity field $u_{0}$ is a vortex, i.e. the stream function is radially symmetric, then the flow evolves but remains a vortex flow for all time (see e.g. [batchelor, Sec. 4.5], [saffman, Chap. 13]). Here, a function $f$ on $\mathbb{R}^{2}$ is said to be radially symmetric if $f(x)$ only depends on $|x|$. The same result holds for the fractional Navier-Stokes equations (Section 5).
@@
$F(f,f)(\xi,t)=\frac{i}{2}\int_{0}^{t}|\xi|e^{-|\xi|^{2\gamma}s}\int_{\mathbb{R}^{2}}\text{sign}(\xi\times\eta)g(\eta,t-s)g(\zeta,t-s)\sin(\theta_{\xi,\eta}-\theta_{\xi,\zeta})d\eta ds$

### v2

@@ 3.3 Numerical results
The quadratic form $a\odot_{\xi}b$ defined by (2.8) satisfies the following estimate:
@@
The proof of (4.1) is elementary and can be found in [18, Lem. 2.1].
To establish the integrability of the solution process $\mathbf{X}$ defined (2.14) it is natural to use the majorization (4.1) of the quadratic form $a\odot_{\xi}b$. In so doing, we establish a natural relation with equation (MS), which will be referred in this paper as the (fractional) Montgomery-Smith equation. The author of [32] introduced (MS) for the case $d=3,\gamma=1$, as a tool to explore the limitation of methods of functional analysis in establishing the well possednes of the Navier-Stokes equations. In the context of $(d,\gamma)$-DSY cascades, the Montgomery-Smith equation plays two roles. The first role is as the governing equation for the probability of no explosion
$\rho(\xi,t)=P_{\xi}(S>t),\quad\rho(\xi,0)=1,$
@@
For the first, recall that in Section 3 of the $(d,\gamma)$-DSY cascade, the conditional distribution of the wave vector $W_{v1}$ given $W_{v}=\xi$ is $H(\eta|\xi)=h(\eta)h(\xi-\eta)/(|\xi|^{2\gamma-1}h(\xi))$ where $h$, given by (2.10), is a positive function satisfying the fundamental relation (2.11) $h*h(\xi)=|\xi|^{2\gamma-1}h(\xi)$. We then have,
@@
Consequently, $\rho$ is continuous in time $t$.
@@
This follows similar lines as the proof of Section 2. Note that $\mathscr{X}(\xi,t)={\mathds{1}_{S_{\xi}>t}}$ satisfies
$\mathscr{X}(\xi,t)=\left\{\begin{array}[]{ll}1,&Y_{\emptyset}\geq t,\\ \mathscr{X}^{(1)}(\eta,t-Y_{\emptyset})\mathscr{X}^{(2)}(\xi-\eta,t-Y_{\emptyset}),\quad&Y_{\emptyset}<t,\end{array}\right$. (4.3)
where $\mathscr{X}^{(1)}$ and $\mathscr{X}^{(2)}$ are two conditionally independent copies on $\mathscr{X}$. Indeed, this follows from the inclusion $Y_{\emptyset}\geq t]\subset[S>t]$ and from the inheritance property of the non-explosion event: ${{\mathds{1}_{S>t}}}{\mathds{1}_{Y_{\emptyset}<t}}={\mathds{1}_{Y_{\emptyset}<t}}\left({\mathds{1}_{{{S}_{{{W}_{1}}}}>t-Y_{\emptyset}}}+{\mathds{1}_{{S_{{{W}_{2}}}}>t-Y_{\emptyset}}}\right)$. Since $|\bar{\mathbf{X}}|\leq 1$, we can take expectation in (4.3) to obtain that $\rho=\mathbb{E}\bar{\mathbf{X}}$ satisfies (4.2). The continuity of $\rho$ follows immediately from (4.2) since $\int_{\mathbb{R}^{3}}\rho(\eta,t-s)\rho(\xi-\eta,t-s)H(\eta|\xi)d\eta\leq 1.$
Note that the constant function $1$ is a trivial solution of (4.2). Consequently, determining whether stochastic explosion occurs, i.e. $\rho(\xi,t)<1$ for some $(\xi,t)$) is equivalent to determining whether (4.2) has other solutions besides the trivial solution for the same initial data $\rho(\xi,0)-1$. This is indeed the case, since as shown in Section 3 there exists values of $(d,\gamma)$ such that the resulting DSY cascade is explosive.
Generalizing the equation introduced in [32], consider the fractional Montgomery-Smith equation in spatial dimension $d\geq 1$ given by
@@
with initial data $v_{0}(\xi)$. Here, $v(\xi,t)$ and $v_{0}(\xi)=v(\xi,0)$ are scalar valued functions. Note that (MS) has the same scaling property as the fractional Navier-Stokes equations (NSE)d,γ. The Fourier transform of (MS) is given, with $c_{0}=(2\pi)^{-d/2}$, by
@@
In what follows it will be necessary to introduce the normalized Fourier Montgomery-Smith equation given by
@@
$\mathbb{E}_{\xi}[|{\bf{X}}|{\mathds{1}_{S>t}}]\leq 2\mathbb{E}_{\xi}\bar{\mathbf{X}}<\infty$
@@
According to Section 4(c), if (FMS) has a mild solution on $(0,\tau)$ then $\mathbb{E}_{\xi}[|{\bf{X}}|\mathbbm{1}_{S>t}]<\infty$ a.e. on $\mathbb{R}^{d}\times(0,\tau)$ which implies the lower bound on the critical time $t_{c}\geq\tau$. The following result (due to Ferreira and Villamizar-Roa [23]) asserts the local-in-time existence and uniqueness of mild solutions to (FMS) for initial data $v_{0}$ in the pseudomeasure-type space $PM^{a}$ defined by (2.3).
@@
Note that the proof of Theorem 1.2, 1.3, and Remark 1.5(b) of [23] for the special case $a=b$ and $\theta=1$ does not use any special structure of the $\odot$-product other than the fact that $|a\odot_{\xi}b|\leq|a||b|$. Thus, the results hold for the fractional Montgomery-Smith equation. ∎
In [18], where the case $d\geq 3$ and $\gamma=1$ is considered, the authors showed that the solution blows up in finite time for sufficiently large compactly supported initial data. The same result holds for the general case $d\geq 1$ and $\gamma\in(\frac{1}{2},\frac{d+2}{4})$. More specifically, we have:
Proposition 4.6.
[…]
Proof.
[…]
$\psi(\xi,t)\geq\sum\limits_{n=1}^{\infty}{{{a}^{n}}{{q}_{n}}(\xi,t)}.$ (4.6)
We have $q_{1}(\xi,t)=e^{-|\xi|^{2\gamma}t}\mathbbm{1}_{D}(\xi)$. By conditioning on the first time of branching, one gets
${{q}_{n}}(\xi,t)=\int_{0}^{t}{|\xi|^{2\gamma}{{e}^{-s|\xi|^{2\gamma}}}\int_{{{\mathbb{R}}^{d}}}{\sum\limits_{k=1}^{n-1}{{{q}_{k}}(\eta,t-s){{q}_{n-k}}(\xi-\eta,t-s)H(\eta|\xi)d\eta ds}}}.$ (4.7)
From now on, the proof follows almost verbatim the proof of Proposition 4.14 in [18], which we will leave to the reader. ∎
Remark 4.7.
The majorizing kernel $h:\mathbb{R}\to(0,\infty]$ does not exist in the case $d=\gamma=1$ (see [34]), so our probabilistic model does not apply to this case. Nevertheless, Sverak and Polacik consider the complex Burgers equation ($d=\gamma=1$) and obtain a similar result [35]. Namely, the equation has a global solution $u\in L^{1}(\mathbb{R}\times(0,1))$ provided that $\|\textup{Im}(u_{0})\|_{L^{1}}\leq 2\pi$. Moreover, they show that the upper bound $2\pi$ is sharp in the sense that it is the supremum of the set of $\rho>0$ such that if $\|u_{0}\|_{L_{1}}<\rho$ then (2.20) is globally well-posed.
@@
We begin by obtaining an expression of the solution operator that involves a product of scalar quantities determined in terms of the initial data, and the angles between the labels of the $(d,\gamma)-DSY$.
@@
Let $N(\xi,t)=|\hskip-3.61371pt\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)|$. We will prove by induction on $N(\xi,t)$. If $N(\xi,t)=0$ then $\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)=\emptyset$, $\partial V(\xi,t)=\{\emptyset\}$ and
@@
Thus, (5.1) holds for the base case $N(\xi,t)=0$. Suppose that (5.1) holds for any $N(\xi,t)\leq n$, for some $n\geq 0$. For $N(\xi,t)=n+1\geq 1$, we have $Y_{\emptyset}<t$. Thus,
$\mathbf{X}(\xi,t)=\mathbf{X}^{(1)}(W_{1},t-Y_{\emptyset})\odot_{\xi}\mathbf{X}^{(2)}(W_{2},t-Y_{\emptyset}).$ (5.2)
@@
$\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)$ $=$ $\{\emptyset\}\cup\stackrel{{\scriptstyle\circ}}{{V}}^{(1)}(W_{1},t-Y_{\emptyset})\cup\stackrel{{\scriptstyle\circ}}{{V}}^{(2)}(W_{2},t-Y_{\emptyset})$ (5.3)
$\partial V(\xi,t)$ $=$ $\partial V^{(1)}(W_{1},t-Y_{\emptyset})\cup\partial V^{(2)}(W_{2},t-Y_{\emptyset})$ (5.4)
where $\stackrel{{\scriptstyle\circ}}{{V}}^{(k)}$ and $\partial V^{(k)}$ are the set of internal nodes and the set of leaves of the subtree rooted at $W_{k}$. Since $N^{(1)}(W_{1},t-Y_{\emptyset})+N^{(2)}(W_{2},t-Y_{\emptyset})=N(\xi,t)-1=n$, one deduces by the induction hypothesis that for $k\in\{1,2\}$,
$\mathbf{X}^{(k)}(W_{k},t-Y_{\emptyset})$ $=$ $\left(\frac{i}{2}\right)^{N^{(k)}(W_{k},t-Y_{\emptyset})}\cdot$
$\left(\prod_{v\in\stackrel{{\scriptstyle\circ}}{{V}}^{(k)}(W_{k},t-Y_{\emptyset})}\textup{sign}(W_{v}\times W_{v1})\sin(\theta_{W_{v},W_{v1}}-\theta_{W_{v},W_{v2}})\right)\cdot$
$\left(\prod_{v\in\partial V^{(k)}(W_{k},t-Y_{\emptyset})}\chi_{0}(W_{v})\cdot e_{W_{v}^{\perp}}\right)e_{W_{k}^{\perp}}$
@@
To use Section 5 to explore the possibilities for stochastic cancellation we consider the well-known simple vortex model, see e.g. [30, p 48]. The stream function corresponding to the initial velocity is radially symmetric, and thus, in this case $\hat{u}_{0}(x)=h(|\xi|)(-\xi_{2},\xi_{1})^{t}$. We will refer to this data as a vortex and note that after scaling, $\chi_{0}(\xi)=\rho(|\xi|)e_{\xi^{\perp}}$ for an appropriate complex valued function $\rho$. Section 5 reveals a stochastic asymmetry that results in substantial cancellations in the proof of Section 5 for this type of initial data.
@@
Let $\tilde{\mathbf{X}}(\xi,t)=\phi_{\xi,t}(\{\tilde{\Sigma}_{v}\}_{v\in\mathbb{T}},\{\tilde{\Theta}_{v}\}_{v\in\mathbb{T}})$ where $\{\tilde{\Sigma}_{v}\}_{v\in\mathbb{T}}$ and $\{\tilde{\Theta}_{v}\}_{v\in\mathbb{T}}$ are given by (5.6)-(5.9). Suppose the vector field $\chi_{0}$ is a vortex. Then
@@
The radial symmetry of the initial data and the invariance of the DSY process leads to the following consequence, also noted in [30, p 48] in the physical space setting.
@@
Suppose $u_{0}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ is a vortex. Then $\textup{(NSE)}_{d,\gamma}$ has a global solution $u$ whose Fourier transform is given by $\hat{u}(\xi,t)=\hat{u}_{0}(\xi)e^{-|\xi|^{2\gamma}t}$.
@@
${\mathbf{X}}^{*}(\xi,t)=\hat{u}_{0}(\xi){\mathds{1}_{[T_{\emptyset}>t]}}.$
@@
Remark 5.5.
Section 5 can also be proven analytically in Fourier space. For convenience, the proof is provided in Section 6. However, our approach highlights the role symmetries of DSY cascades play in providing cancellations necessary to prove global existence of solutions of $\textup{(NSE)}_{d,\gamma}$.
@@
For each $r>0$, let $D_{r}=\{\xi\in\mathbb{R}^{2}:|\xi|<r\}$. By the continuity of the translation operator in $L^{1}(\mathbb{R}^{2})$ (see e.g. [36, Exer. 12, p. 132]),
@@
Thus, $q(\tau)\geq\limsup_{t\to\tau^{-}}p(t)=\infty$. This is a contradiction. ∎ The fact that the solution process $\mathbf{X}(\xi,t)$ fails to be integrable after finite time for sufficiently large initial condition does not necessarily imply that the mild solution $u$ to $\textup{(NSE)}_{d,\gamma}$ blows up after finite time. It simply implies that the mean-flow representation of the scaled Fourier transform $\chi=c_{0}\frac{\hat{u}}{h}$ ceases to be valid after finite time. For example, if the initial velocity field $u_{0}$ is a vortex, i.e. the stream function is radially symmetric, then the flow evolves but remains a vortex flow (radial eddy) for all time. See e.g. [30, Example 2.2, p. 48], [5, Sec. 4.5], [37, Chap. 13]. Here, a function $f$ on $\mathbb{R}^{2}$ is said to be radially symmetric if $f(x)$ only depends on $|x|$. The same result holds for the fractional Navier-Stokes equations (Section 6).
@@ 6 Appendix: Analytical Proof of Section 5
$F_{2}(f,f)(\xi,t)=\frac{i}{2}\int_{0}^{t}|\xi|e^{-|\xi|^{2\gamma}s}\int_{\mathbb{R}^{2}}\text{sign}(\xi\times\eta)g(\eta,t-s)g(\zeta,t-s)\sin(\theta_{\xi,\eta}-\theta_{\xi,\zeta})d\eta ds$